OpenStudy (anonymous):

Experimental or Theoretical Probability? You toss a coin 50 times and you end up with 20 heads and 30 tails P(H) = 20/50 = 0.4 P(T) = 30/50 = 0.6 Experimental Theoretical Cannot be determined Neither

OpenStudy (amistre64):

are the results an true reflection of an experiment? or are they approximations and never actually happened in the experiment?

OpenStudy (anonymous):

so im guessing its experimental @amistre64

OpenStudy (amistre64):

experimental is a good guess :)

OpenStudy (amistre64):

we can assess the 'true' probability if need be by saying: \[P(H)=.4\pm 1.96(\sqrt{\frac{.4(.6)}{50}})\] if memory serves, 1.96 is the zscore related to 95% of the data about the mean.

OpenStudy (anonymous):

oh that means its theoretical @amistre64

OpenStudy (amistre64):

yeah, if we have to adjust stuff to get it to work in out theory, then its theoretical if we just get the results from performing some experiment, then its experimental.

OpenStudy (anonymous):

yup so its theo ratical

OpenStudy (amistre64):

we got 20 heads out of 50 tosses P(H) = 20/50 is directly from the experiment, so it has to be experimental